# Learning System Dynamics without Forgetting

### v1

@@ 3.3 Masked Encoder Network
As the first component of the model, the masked encoder network serves to encode the dynamic pattern in the observational data into the latent space. As introduced in Section 3.1, the observation of the system is a spatial temporal graph. Therefore, we adopt the attention-based spatial temporal graph neural network framework (ST-GNN) [38, 7, 8, 39] to encode the observational data. Originally, the graph attention network (GAT) [40] is designed to aggregate information over the spatially neighboring nodes. On spatial temporal graphs, the information aggregation is extended to both spatial and temporal neighboring nodes. Such a spatial temporal neighborhood of a state $\mathbf{x}_{v}^{t}$ is defined as the states of the spatially connected nodes within a specified temporal window $\delta_{window}$,
@@
Then, the information aggregation is conducted over the spatial temporal neighborhood to iteratively update the representation of each state. Formally, the update of the hidden representation of a state $\mathbf{x}_{v}^{t}$ at the $l$-the layer of the ST-GNN is formulated as,
@@
where $\mathrm{a}(\cdot,\cdot)$ calculates the attention score between a pair of states, and $\mathrm{msg}(\cdot,\cdot)$ is the message function commonly adopted in graph neural networks [41]. However, different from the message function in typical graph neural networks, the temporal relationship between the states is a crucial part to understand the system dynamics in our task. Therefore, we follow the strategy in LG-ODE [7] to incorporate the temporal distance between the states in the message function and attention function,
@@
where $\mathrm{TE(\cdot)}$ is a temporal position encoding developed based on the position encoding in Transformer [42] for incorporating the temporal information into the representation. Finally, for the subsequent state prediction task, the representations of the states are averaged over the temporal dimension,
@@
where $t_{0}$ denotes the starting time of all states in the observational data (Section 3.1), and the average term $\bar{\mathbf{h}}_{v}$ of each node $v$ is calculated by weighted summation over the representations at all time steps,
@@ 3.4 Masked ODE-based Generator
For practical consideration, we develop the masked ODE-based generator based on the latent ODE framework [43, 7, 21], such that the model can handle observation with irregular temporal intervals and predict future states at any time denoted by real numbers. Specifically, the trajectory prediction problem is formulated as as solving an ODE initial value problem (IVP), where the initial values of the objects ($\{z_{v}^{t_{1}}|v\in\mathbb{V}\}$) are generated from the final representation of the observational data ($\{\mathbf{h}_{final}^{v}|v\in\mathbb{V}\}$, Section 3.3). Mathematically, the procedure of predicting the future trajectory of a system $s$ is formulated as,
@@
As noted at the end of Section 3.2, the parameters of $\mathrm{mlp}(\cdot)$ and $\mathrm{F}_{int}(\cdot)$ are part of $\theta_{G}$ and controlled by $\mathbf{M}_{G}^{s}$.
Based on the approximate posterior distribution $q(\{z_{v}^{t_{1}}|v\in\mathbb{V}\}|\{\mathbb{X}_{v}|v\in\mathbb{V}\})$, we sample an initial state for each object, upon which the ODE solver will be applied for generating the predicted states in the latent space. The dynamics of each object in the system is governed by its interaction with all the other objects. Therefore, the core part of the ODE-based generator is a trainable interaction network that encodes the dynamics in the form of the derivative of each $\mathbf{z}_{v}$,
@@
where the function $\mathrm{F}_{int}(\cdot)$ parameterized by $\theta_{I}$ predicts the dynamics (derivative) of each object $v$ based on all the other objects interacting with $v$ (i.e. $\mathcal{N}_{s}(v)$ defined in Section 3.1), and $t^{\prime}$ denotes any possible future time. Note that $\mathcal{N}_{s}(v)$ only contains the spatial neighbors and is different from $\mathcal{N}_{st}(\cdot)$. Because the object-wise interaction at a certain time $t^{\prime}$ is not dependent on system states at other times. For example, in a charged particle system governed by electrostatic force, the force between each pair of particle at $t^{\prime}$ is solely determined by the relative positions (i.e. the states) of the particles (Figure 2) at $t^{\prime}$. In our work, we adopt the Neural Relational Inference (NRI) [3] as the function $\mathrm{F}_{int}(\cdot)$. Based on $\mathrm{F}_{int}(\cdot)$, the future state of the system at an arbitrary future time $t_{2}$ can be obtained via an integration
@@ 3.5 Mask Optimization & Selection
where $p(\mathbf{Z}_{\mathbb{V}}^{t_{1}})$ denotes the prior distribution of $\mathbf{Z}_{\mathbb{V}}^{t_{1}}$, which is typically chosen as standard Gaussian. $\mathbf{M}^{s}$ denotes the union of all masks over different modules of the framework (i.e. $\mathbf{M}^{s}=\{\mathbf{M}_{E}^{s},\mathbf{M}_{I}^{s},\mathbf{M}_{P}^{s}\}$). In our model, the parameters ($\theta_{E}$,$\theta_{I}$, and $\theta_{P}$) will be fixed, and the ELBO will be maximized by optimizing the binary masks $\mathbf{M}^{s}$ overlaying the parameters via the Edge-popup algorithm [11]. After learning each system, the obtained mask is added into a mask pool $\mathbb{M}$ to be used in testing. During testing, MS-GODE will automatically select the most appropriate mask based on the confidence of reconstructing part of the given observations. Specifically, a given observations from [$t_{0}$,$t_{1}$] is first split it into two periods [$t_{0}$,$\frac{t_{0}+t_{1}}{2}$] and [$\frac{t_{0}+t_{1}}{2}$,$t_{1}$]. Then, the first half is fed into the model and the correct mask can be chosen by selecting the one that can reconstruct the second half with the lowest error.
@@ 4 Experiments
In this section, we aim to answer the following questions. 1. How to properly configure MS-GODE for optimal performance. 2. How would the configuration of the system sequence influence the performance? 3. How is the performance of the existing CL techniques? 4. Can MS-GODE outperform the baselines?
@@ 4.1 Experimental Systems
In experiments, we adopt physics and biological cellular systems. The detailed introduction on the system sequence construction is provided in Appendix A.1.
Simulated physics systems are commonly adopted to evaluate the machine learning models in the task of learning system dynamics [6, 2, 7]. The physics systems adopted in this work include spring connected particles and charged particles (Figure 2). We carefully adjust the system configuration and construct 3 system sequences with different level of dynamics shift (Appendix A.1).
Biological cellular systems are innovatively introduced in this work based on Virtual Cell platform [13, 14, 15]. In experiments, we adopt two models: EGFR receptor interaction model and Ran protein based translocation model. We adjust the coefficients of the models to construct a sequence containing 2 EGFR and 2 Ran systems interleaved with each other ($\mathrm{EGFR}_{1}\rightarrow\mathrm{Ran}_{1}\rightarrow\mathrm{EGFR}_{2}\rightarrow\mathrm{Ran}_{2}$). Full details are provided in Appendix A.1.2.
@@ 4.2 Experimental Setups & Evaluation
Model evaluation under the continual learning setting. Different from standard learning setting, the models in this paper learns sequentially on multiple systems under the continual learning and evaluation are significantly different. After learning each new task, the model is tested on all learned tasks and the results form a performance matrix $\mathrm{M}^{p}\in\mathbb{R}^{N\times N}$, where $\mathrm{M}^{p}_{i,j}$ denotes the performance on the $j$-system after learning from the $1$-st to the $i$-th system, and $N$ is the number of systems in the sequence. In our experiments, each entry of $\mathrm{M}^{p}$ is a mean square error (MSE) evaluating the performance on a single system. To evaluate the performance over all systems, average performance (AM) can be calculated. For example, $\frac{\sum_{j=1}^{i}\mathrm{M}^{p}_{N,j}}{N}$ is the average performance after learning the entire sequence with $N$ tasks. Similarly, average forgetting (AF) can be calculated as $\frac{\sum_{j=1}^{N-1}\mathrm{M}^{p}_{N,j}-\mathrm{M}^{p}_{j,j}}{N-1}$. More details on model evaluation can be found in Appendix A.4. All experiments are repeated 5 times on a Nvidia Titan Xp GPU. The results are reported with average and standard deviations.
Baselines & model settings. We adopt state-of-the-arts baselines including the performance upper (joint training) and lower bounds (fine-tune) in experiments. The state-of-the-arts baselines adopted in this work include Elastic Weight Consolidation (EWC) [28] based on regularization, Learning without Forgetting (LwF) [44] based on knowledge distillation, Gradient Episodic Memory (GEM) [34] based on both memory replay and regularization, Bias-Correction [45] based Memory Replay (BCMR) that integrates the idea of data sampling bias correction into memory replay, and Scheduled Data Prior (SDP) [46] that adopts a data-driven approach to balance the contribution of past and current data. Besides, joint training and fine-tune are commonly adopted in continual learning works. Joint training trains the model jointly over all systems, which does not follow the continual learning setting. Fine-tune directly trains the model incrementally on new systems without any continual learning technique.
More details of experimental settings and baselines and experimental settings are provided in Appendix A.1 A.3.
@@ 4.3 Model Configuration and Performance (RQ1)
When optimizing the system-specific masks using the edge-popup algorithm [11] (Appendix A.2), each entry of the mask is assigned with a continuous value for gradient descent after backpropagation. During inference, different strategies can be adopted to transform the continuous scores into binary values. In our experiments, we tested both ‘fast selection’ and ‘top-k selection’ with different thresholds. ‘Fast selection’ set all entries with positive score values into ‘1’s and the other entries into ‘0’s. ‘Top-k selection’ first ranks the score values and set a specified ratio of entries with the largest values into ‘1’s. In Figure 4, we show the performance of different strategies. Overall, ‘top-k selection’ is inferior to ‘fast selection’. This is potentially because that ‘fast selection’ does not limit the number of selected entries, therefore allowing more flexibility for optimization. We also observe that the performance of ‘top-k’ selection is more sensitive on the cellular systems compared to the physics systems, indicating that the cellular systems have higher optimization difficulty for sub-network (binary mask) learning over system sequences.
Second, sub-network learning will deactivate some neurons in the model, which resembles the dropout mechanism widely adopted in machine learning models and may cause the model to be over sparsified. Therefore, we investigate the influence of dropout rate in MS-GODE. From Figure 5, we could see that smaller dropout rate results in lower error (better performance). This corroborates our hypothesis that the mask selection mechanism and dropout are complementing each other, and the dropout rate should be decreased when the masking strategy is adopted.
@@
4.4 Sequence Configuration and Performance (RQ2)
In this subsection, we investigate the influence of system sequence configuration on the learning difficulty and model performance. Specifically, we construct 3 physics system sequences with increasing level of dynamics shift (Details are provided in Appendix A.1.1). Sequence 1 (low-level dynamics shift) consists of 8 spring connected particle systems, in which consecutive systems are only different in one system coefficient. Sequence 2 (mid-level dynamics shift) is constructed to have higher level of dynamics shift by simultaneously varying 2 system coefficients. Finally, sequence 3 (high-level dynamics shift) is constructed by interleaving spring connected particle systems and charged particle systems with disparate dynamics. As shown in Table 1, in terms of both AP and AF, most methods, including MS-GODE, obtain similar performance over Sequence 1 and 2, and obtain better performance on sequence 3. This is reasonable for MS-GODE, since the systems with more diverse dynamics are easier to distinguish for selecting the masks during inference. For the other baselines, e.g. Fine-tune, the results indicate that diverse systems may guide the model to exploit different parameters for different systems, while similar systems may cause the same part of parameters to be updated more frequently, resulting in more severe forgetting issue.

### v2

@@ 3.3 Masked Encoder Network
As the first component of the model, the masked encoder network serves to encode the dynamics pattern in the observational data into the latent space. As introduced in Section 3.1, the trajectory of a system is a spatial-temporal graph. Therefore, the encoder is constructed as an attention-based spatial-temporal graph neural network framework (ST-GNN) (Huang et al., 2020; Hu et al., 2022; Huang et al., 2021; Zhang et al., 2020a). Originally, the graph attention network (GAT) (Veličković et al., 2017) is designed to aggregate information over the spatially neighboring nodes. On spatial-temporal graphs, the information aggregation is extended to both spatial and temporal neighboring nodes. Such a spatial-temporal neighborhood of a state $\mathbf{x}_{v}^{t}$ is defined as the states of the spatially connected nodes within a specified temporal window $\delta_{window}$,
@@
Then, iterative message passing is conducted over the spatial-temporal edges to update the representation of each state. The update of the hidden representation of a state $\mathbf{x}_{v}^{t}$ at the $l$-the layer of the ST-GNN is formulated as,
@@
where $\mathrm{a}(\cdot,\cdot)$ calculates the attention score between a pair of states, and $\mathrm{msg}(\cdot,\cdot)$ is the message function commonly adopted in GNNs (Gilmer et al., 2017). However, different from the message function in typical GNNs, the temporal relationship between the states is a crucial part to understand the system dynamics. Therefore, we follow the strategy in LG-ODE (Huang et al., 2020) to incorporate the temporal distance between the states in the message function and attention function,
@@
where $\mathrm{TE(\cdot)}$ is a temporal position encoding developed based on the position encoding in Transformer (Vaswani et al., 2017) for incorporating the temporal information into the representation. Finally, for subsequent state prediction, the state representations are averaged over the temporal dimension,
@@
where $t_{0}$ denotes the starting time of all states in the observational data (Section 3.1), and the average term $\bar{\mathbf{h}}_{v}$ of each node $v$ is a weighted summation over the representations at all time steps,
@@ 3.4 Masked ODE-based Generator
ODE-based generator (Huang et al., 2020; Chen et al., 2018; Rubanova et al., 2019) ensures that the model can handle observations with irregular temporal intervals and incomplete states, as well as predict future states at any time denoted by real numbers. Specifically, the trajectory prediction is formulated as solving an ODE initial value problem (IVP), where the initial values of the objects ($\{z_{v}^{t_{1}}|v\in\mathbb{V}\}$) are generated from the final representation of the states ($\{\mathbf{h}_{final}^{v}|v\in\mathbb{V}\}$, Section 3.3). Mathematically, the procedure of predicting the future trajectory of a system $s$ is formulated as,
@@
As noted in Section 3.2, the parameters of $\mathrm{mlp}(\cdot)$ and $\mathrm{F}_{int}(\cdot)$ are part of $\theta_{G}$ and controlled by $\mathbf{M}_{G}^{s}$.
Based on the approximate posterior distribution $q(\{z_{v}^{t_{1}}|v\in\mathbb{V}\}|\{\mathbb{X}_{v}|v\in\mathbb{V}\})$, we sample an initial state for each object, upon which the ODE solver will be applied for generating the predicted states in the latent space. The dynamics of each object in the system are governed by its interaction with all the other objects. Therefore, the core part of the ODE-based generator is a trainable interaction network that encodes the dynamics in the form of the derivative of each $\mathbf{z}_{v}$,
@@
where the function $\mathrm{F}_{int}(\cdot)$ parameterized by $\theta_{I}$ predicts the dynamics (derivative) of each object $v$ based on all the other objects interacting with $v$ (i.e. $\mathcal{N}_{s}(v)$ defined in Section 3.1), and $t^{\prime}$ denotes any possible future time. Note that $\mathcal{N}_{s}(v)$ only contains the spatial neighbors and is different from $\mathcal{N}_{st}(\cdot)$, because the object-wise interaction at a certain time $t^{\prime}$ is not dependent on system states at other times. For example, in a charged particle system governed by electrostatic force, the force between a pair of particles at $t^{\prime}$ is solely determined by the relative positions (i.e. the states) of the particles (Figure 7) at $t^{\prime}$. In our work, we adopt the Neural Relational Inference (NRI) (Kipf et al., 2018) as the function $\mathrm{F}_{int}(\cdot)$. Based on $\mathrm{F}_{int}(\cdot)$, the future state of the system at an arbitrary future time $t_{2}$ can be obtained via an integration
@@ 3.5 Sub-network Learning
where $p(\mathbf{Z}_{\mathbb{V}}^{t_{1}})$ denotes the prior distribution of $\mathbf{Z}_{\mathbb{V}}^{t_{1}}$, which is typically chosen as standard Gaussian. $\mathbf{M}^{s}$ denotes the union of all masks over different modules of the framework (i.e. $\mathbf{M}^{s}=\{\mathbf{M}_{E}^{s},\mathbf{M}_{I}^{s},\mathbf{M}_{P}^{s}\}$). In our model, the parameters ($\theta_{E}$,$\theta_{I}$, and $\theta_{P}$) will be fixed, and the ELBO will be maximized by optimizing the binary masks $\mathbf{M}^{s}$ overlaying the parameters via the Edge-popup algorithm (Ramanujan et al., 2020). After learning each system, the obtained mask is added into a mask pool $\mathbb{M}$ to be used in testing.
3.6 Mode-Switching
During testing, MS-GODE will be evaluated on a sequence of systems exhibiting diverse dynamics patterns, and it will automatically switch to the optimal mode that ensures the highest accuracy. This is achieved by applying the most suitable mask based on the performance of reconstructing part of the given observations. To obtain the most suitable mask, a given observation from [$t_{0}$,$t_{1}$] is first split it into two periods [$t_{0}$,$\frac{t_{0}+t_{1}}{2}$] and [$\frac{t_{0}+t_{1}}{2}$,$t_{1}$]. Then, the first half is fed into the model and the correct mask can be chosen by selecting the one that can reconstruct the second half with the lowest error.
@@ 4 Experiments
In this section, we aim to answer the following questions. 1. How to properly configure MS-GODE for optimal performance? 2. How would the configuration of the system sequence influence the performance? 3. How is the performance of the existing CL techniques? 4. Can MS-GODE outperform the baselines?
@@ 4.1 Experimental Systems
In experiments, we adopt physics and biological cellular systems. A detailed introduction to the system sequence construction is provided in Appendix A.1.
Simulated physics systems are commonly adopted to evaluate the machine learning models in the task of learning system dynamics (Liu et al., 2024; Battaglia et al., 2016; Huang et al., 2020). The physics systems adopted in this work include spring-connected particles and charged particles (Figure 7). We carefully adjust the system configuration and construct 3 system sequences with different levels of dynamics shift (Appendix A.1).
Biological cellular systems are innovatively introduced in this work based on Virtual Cell platform (Schaff et al., 1997; Cowan et al., 2012; Blinov et al., 2017). Currently, Bio-CDL includes rule-based model of EGFR and compartmental rule based model of translocation based on the Ran protein (GTPase). In experiments, we adjust the coefficients of the models to construct a sequence containing 2 EGFR and 2 Ran systems interleaved with each other ($\mathrm{EGFR}_{1}\rightarrow\mathrm{Ran}_{1}\rightarrow\mathrm{EGFR}_{2}\rightarrow\mathrm{Ran}_{2}$). Full details are provided in Appendix A.1.2.
@@ 4.2 Experimental Setups & Evaluation
Model evaluation in CDL. The models in this work learn sequentially on multiple systems under the continual learning setting, thus the evaluation is significantly different from standard learning settings. After learning each new task, the model is tested on all learned tasks and the results form a performance matrix $\mathrm{M}^{p}\in\mathbb{R}^{N\times N}$, where $\mathrm{M}^{p}_{i,j}$ denotes the performance on the $j$-system after learning from the $1$-st to the $i$-th system, and $N$ is the number of systems in the sequence. In our experiments, each entry of $\mathrm{M}^{p}$ is a mean square error (MSE) evaluating the performance on a single system. To evaluate the performance over all systems, average performance (AP) can be calculated. For example, $\frac{\sum_{j=1}^{N}\mathrm{M}^{p}_{N,j}}{N}$ is the average performance after learning the entire sequence with $N$ tasks. Similarly, average forgetting (AF) can be calculated as $\frac{\sum_{j=1}^{N-1}\mathrm{M}^{p}_{N,j}-\mathrm{M}^{p}_{j,j}}{N-1}$. More details on model evaluation can be found in Appendix A.4. The baseline are configured by combining existing continual learning techniques with the LG-ODE (Huang et al., 2020) model. Other dynamics learning models were excluded because they typically don’t support the practical setting with irregular and incomplete observation studied in this work. All experiments are repeated 5 times on a Nvidia Titan Xp GPU. The results are reported with average and standard deviations.
Baselines & model settings. We adopt state-of-the-art baselines including the performance upper (joint training) and lower bounds (fine-tune) in experiments. The state-of-the-art baselines adopted in this work include Elastic Weight Consolidation (EWC) (Kirkpatrick et al., 2017) based on regularization, Learning without Forgetting (LwF) (Li & Hoiem, 2017) based on knowledge distillation, Gradient Episodic Memory (GEM) (Lopez-Paz & Ranzato, 2017) based on both memory replay and regularization, Bias-Correction (Chrysakis & Moens, 2023) based Memory Replay (BCMR) that integrates the idea of data sampling bias correction into memory replay, and Scheduled Data Prior (SDP) (Koh et al., 2023) that adopts a data-driven approach to balance the contribution of past and current data. Besides, joint training and fine-tuning are commonly adopted in continual learning works. Joint training trains the model jointly over all systems, which does not follow the continual learning setting. Fine-tune directly trains the model incrementally on new systems without any continual learning technique. As revealed by Ramanujan et al. (2020), for learning subnetworks, pre-trained models are the most suitable for serving as the backbone. However, unlike the Computer Vision (CV) tasks studied by Ramanujan et al. (2020), pre-trained models are not readily available in the context of dynamics system modeling. For example, Seifner et al. (2024) provides a promising pre-trained model with valuable insights into pre-training for dynamical systems, but the model is still limited to interpolation tasks. However, we will investigate how to adapt the provided model to extrapolation tasks once their code and model are released. Therefore, we adopt the random initialization strategy, which is shown by Ramanujan et al. (2020) to be less effective but could be comparable to pre-trained models with a proper ratio of remaining parameters, which is thoroughly investigated in our experiments reported in Section 4.3 (Figure 4). More details of experimental settings and baselines and experimental settings are provided in Appendix A.1 A.3.
@@ 4.3 Model Configuration and Performance (RQ1)
When optimizing the system-specific masks using the edge-popup algorithm (Ramanujan et al., 2020) (Appendix A.2), each entry of the mask is assigned with a continuous value for gradient descent after backpropagation. During inference, different strategies can be adopted to transform the continuous scores into binary values. In our experiments, we tested both ‘fast selection’ and ‘top-k selection’ with different thresholds. ‘Fast selection’ sets all entries with positive score values into ‘1’s and the other entries into ‘0’s. ‘Top-k selection’ first ranks the score values and sets a specified ratio of entries with the largest values into ‘1’s. In Figure 4, we show the performance of different strategies. Overall, ‘top-k selection’ is inferior to ‘fast selection’. This is potentially because ‘fast selection’ does not limit the number of selected entries, therefore allowing more flexibility for optimization. We also observe that the performance of ‘top-k’ selection is more sensitive on the cellular systems compared to the physics systems, indicating that the cellular systems have higher optimization difficulty for sub-network (binary mask) learning over system sequences.
Second, sub-network learning will deactivate some neurons in the model, which resembles the dropout mechanism widely adopted in machine learning models and may cause the model to be over-sparsified. Therefore, we investigate the influence of dropout rate in MS-GODE. From Figure 5, we can see that a smaller dropout rate results in lower error (better performance). This corroborates our hypothesis that the mask selection mechanism and dropout complement each other, and the dropout rate should be decreased when the masking strategy is adopted.
@@
4.4 Sequence Configuration and Performance (RQ2)
In this subsection, we investigate the influence of system sequence configuration on the learning difficulty and performance. Specifically, we construct 3 physics system sequences with increasing level of dynamics shift (Appendix A.1.1). Sequence 1 (low-level dynamics shift) consists of 8 spring connected particle systems, in which consecutive systems are only different in one system coefficient. Sequence 2 (mid-level dynamics shift) is constructed to have a higher level of dynamics shift by simultaneously varying 2 system coefficients. Finally, sequence 3 (high-level dynamics shift) is constructed by interleaving spring-connected particle systems and charged particle systems with disparate dynamics. As shown in Table 1, in terms of both AP and AF, most methods, including MS-GODE, obtain similar performance over Sequence 1 and 2, and obtain better performance on Sequence 3. For MS-GODE, since the systems with more diverse dynamics are easier to distinguish for selecting the masks during inference. For all the methods in general, two possible cases are: 1. The model is not well trained on any system (it has not encoded much information), therefore it has nothing to forget. 2. The model is well trained and can maintain enough information from each system. However, the lower AP (low MSE) obtained on the system sequence with high-level dynamics shift indicates that the model has well adapted to each system, i,e, case 1 is not true. When case 2 holds, it indicates that the model’s parameters are sufficiently updated to fit each system. When learning on one system, two possible cases are: 1. All parameters are modified uniformly. 2. A subset of parameters are modified more than the others. When the baseline is Fine-tune, if all the parameters are uniformly modified, it would be almost impossible to maintain the performance on previous systems. Therefore, the potential case is that each system relies more on a subset of the parameters, and systems with larger difference in dynamics may lead to less overlap in the subsets of parameters they rely on.